What is the frequency of radiation with a wavelength of 3.82 x 10-7 m?


1
Expert's answer
2021-03-02T04:55:37-0500

The wavelength of radiation is "λ = 3.82 \times 10^{-7} \;m"

The speed of radiation is "c = 3 \times 10^8 \; m/s"

The frequency v of radiation:

"v = \frac{c}{λ} \\

= \frac{3 \times 10^8}{3.82 \times 10^{-7}} \\

= 7.85 \times 10^{14} \;s^{-1} \\

= 7.85 \times 10^{14} \;Hz"